### v4

# WakenLLM: Evaluating Reasoning Potential and Stability in LLMs via Fine-Grained Benchmarking

Abstract
Large Language Models (LLMs) frequently output the label Unknown in reasoning tasks, where two scenarios may appear: (i) an input sample is genuinely unverifiable, but the model cannot understand why; and (ii) a verifiable problem that the model fails to solve, and thus the model outputs Unknown. We refer to these cases collectively as the Vague Perception phenomenon. Current evaluations focus on whether such answers are honest, rather than analyzing the limits of LLM reasoning. To address this, we introduce WakenLLM, a framework that quantifies the portion of Unknown output attributable to model incapacity and evaluates whether stimulation can convert them into either correct answers (verifiable samples) or justified explanations (unverifiable samples) supported by valid reasoning. Our method offers a clearer picture of the limits of LLM reasoning and the potential for corrections across various datasets. Comprehensive experiments on six LLMs suggest that, without further training or parameter revision, LLMs can achieve up to a 68.53% increase in accuracy on Vague Perception samples through guided understanding. Our work reveals that current baseline methods only activate a small portion of LLMs’ reasoning potential, indicating considerable unexplored capacity. This extends the empirical upper bounds of reasoning accuracy in LLMs. Consequently, this study deepens our understanding of the latent reasoning capacity of LLMs and offers a new perspective on addressing the Vague Perception phenomenon.
WakenLLM: Evaluating Reasoning Potential and Stability in LLMs via Fine-Grained Benchmarking
Zipeng Ling2∗, Yuehao Tang1∗, Shuliangliu1∗, Junqi Yang3∗, Shenghong Fu5 Chen Huang3, Kejia Huang4, Yao Wan3, Zhichao Hou2, Xuming Hu1 1Hong Kong University of Science and Technology (Guangzhou), 2University of Pennsylvania 3Huazhong University of Science and Technology 4Nanjing University of Posts and Telecommunications, 5Hong Kong Polytechnic University {zpling0816, xuminghu97}@gmail.com
Figure 1: Definition of Vague Perception: when an LLM misjudges a verifiable sample as Unknown, or it is unable to tell why a sample should be categorized as unknown when facing objectively unsolvable ones. The figure presents two representative reasons attributed to this phenomenon: Fact Understanding and Reasoning Gap—highlighting 1. Limited ability to solve a problem. 2. Misunderstandings of facts versus failures to link them.
1 Introduction
Large Language Models (LLMs) such as GPT‑3, PaLM, and GPT‑4 have shown impressive zero‑shot and few‑shot reasoning on benchmarks like BIG‑Bench, MMLU, and GSM8K Brown et al. (2020); OpenAI (2023); Chowdhery et al. (2022); Srivastava and et al. (2022); Hendrycks and et al. (2021); Cobbe et al. (2021). Their performance is further improved by prompting techniques such as Chain-of-Thought, Self‑Consistency, and Program‑Aided Reasoning Kojima et al. (2022); Wei et al. (2022b); Wang et al. (2023c); Lewkowycz and et al. (2022). Meanwhile, most evaluations rank the overall trustworthiness the first and causal atmosphere secondJiang et al. (2024); Bubeck (2023). Recent diagnostic benchmarks—such as potential‑space reasoning Hagendorff and Fabi (2025), visual hallucination stability Liu et al. (2025), and metaphysical reasoning (MARS) Wang and Song (2024)—demonstrate the need for more fine‑grained evaluations.
Table 1: Comparison of different LLM reasoning–guidance frameworks. WakenLLM integrates all eight dimensions as the first holistic solution for diagnosing and mitigating the Vague Perception phenomenon. For used symbols, : Fully covered, : Partially covered, : Not covered.
Paradigm Guidance Questioning Stimulation
Dynamic Samples
Selection
Reflection
Stage-wise
Examining
Fine-Grained
Decomposing
Diversity
CoT Wei et al. (2022b)
ToT Yao et al. (2023b)
Faithful-CoT Lyu et al. (2023)
Plan&Solve Wang et al. (2023b)
Self-Refine Madaan et al. (2023)
Reflexion Shinn et al. (2023)
Decomposed Prompting Khot et al. (2023)
Active Prompt Diao et al. (2023)
SCORE Zhang et al. (2024)
ProCo Wu et al. (2024)
CoTAL Cohn et al. (2025)
Cochain Zhao et al. (2025)
ISP2 Zhu et al. (2025)
Trace-of-Thought McDonald and Emami (2025)
MAPS de Souza Loureiro et al. (2025)
CoD Xu et al. (2025)
ReSearch Chen et al. (2025)
WakenLLM (Ours)
Fine-grained reasoning reliability remains under-explored. Empirical studies have shown that LLMs frequently hallucinate, or output “Unknown” even when the context contains sufficient evidence, while at other times they output Unknown without understanding why, facing fundamentally under-specified Lin (2023); Zhou et al. (2025) problems. This paper formalizes this phenomenon as Vague Perception: a mismatch between a model’s reasoning ability and the objective problems. Therefore, a critical gap lies in understanding how LLMs deal with complex reasoning and uncertainty.
In this work, we analyze the root causes of deep-reasoning failures behind Vague Perception and propose fine-grained mitigation strategies. To enable systematic measurement, we introduce a new benchmark that spans four subject domains and multiple subcategories, offered in two different formats: (i) Fact-based benchmark, where relevant information is listed as discrete atomic statements (e.g., “A is X, B is Y”); and (ii) Story-based benchmark, where the same facts are embedded in a coherent passage. This design probes how presentation style and information density influence the emergence and eventual resolution of Vague Perception phenomenon.
We construct benchmarks in which half of the samples are verifiable samples with T/F labels and the other half are unverifiable samples with Unknown labels that are objectively unsolvable. These samples are then fed into the model to extract the corresponding Vague Perception samples. As shown in Figure 1, LLMs sometimes output “Unknown” without a clear understanding of the sample. Subsequently, the selected ones are processed via a two-stage stimulation mechanism aimed at “waking up” the model’s reasoning abilities. By measuring “Awaken Ability", we present a novel perspective on evaluating LLM reasoning, one that considers not only the model’s initial accuracy, but also its potential to be guided for correction.
We propose WakenLLM as a strong baseline to explore the reasoning abilities of LLMs among Vague Perception samples, yielding comparable performance across six different LLMs in these reasoning tasks, highlighting that both open-source and closed-source LLMs show substantial improvements in reasoning accuracy.
To summarize, our contributions are threefold:
  • We dive into Vague Perception, a phenomenon in reasoning tasks that models output unknown because of subjective incapacity. To examine it comprehensively, we propose WakenLLM, a unified framework that quantifies LLMs’ capacity to convert to correct answers with stimulation—offering a principled solution for handling Vague Perception in reasoning tasks.
  • Through systematic analyses across LLMs of varying parameter scales, we devise benchmark metrics that more precisely characterize the challenges posed by Vague Perception. The resulting insights not only boost reasoning accuracy but also deepen our understanding of LLM trustworthiness and advanced reasoning paradigms.
  • By comparing WakenLLM with diverse baselines, we measure how effectively each approach unlocks LLMs’ potential. Incorporating both training and non-training frameworks, our work offers a new perspective on mitigating the Vague Perception phenomenon.
Figure 2: Overview of the proposed pipeline. We begin with a balanced Fact and Story-based dataset, sampling from verifiable (T/F) and unverifiable (Unknown) samples at 1 : 1 ratio. The pipeline then (1) detects cases that expose the model’s Vague Perception during reasoning and (2) applies a two-stage remedy: Stage 1 uses stimulation to re-evaluate the mislabeled samples, while Stage 2 gathers the remaining False Converting samples from Stage 1 for model self-reflection. Stimulation Types vary with data form.
2 Evaluation Framework
The WakenLLM evaluation framework, illustrated in Figure 2, is composed of two successive stages. It evaluates the model’s latent reasoning ability and systematically evaluates the challenge of Vague Perception. For the vanilla setting, the first stage assesses the initial Vague Perception samples, while the second investigates the false converting samples $FC^{2}$, allowing us to compute the Overall Benchmark Converting Rate, indicating the percentage of Vague Perception that can be correctly reasoned by LLMs. Under the Remind-then-Guide (RtG) setting, in addition to reminding the model of its prior mistake, we either assign labels—true or false with a preset probability—or provide the previous reasoning process as input to examine whether the LLM still exhibits conformity. We examine whether LLMs still show conformity after being reminded that they solved the problems wrong previously, or it would learn from it and distinguish whether it is guiding or misguiding.
2.1 Overall Benchmark Converting Rate
Our framework begins with a model-specific pre-processing stage that identifies Vague Perception samples—cases where the LLM initially mislabels (e.g., saying Unknown when the sample is verifiable, or failing to explain an unverifiable sample). Since this process is performed separately for each model, the set of Vague Perception samples differs across each model.
After feeding the Stage 1 pipeline with Vague Perception samples, the conversion results fall into three categories: (1) The model fails to generate a coherent logic chain and remains output Unknown, which we call Unexcited Converting (UC). (2) The model produces a false logic chain and an incorrect final label, denoted as False Converting (FC). (3) The model outputs a correct label under stimulation, called True Converting (TC). To quantify these proportions, we use UCR, FCR, and TCR to represent the rates of each type of conversion. Given our focus on the reasoning correction of the model, we elaborate only on TCR and OCR, defining True Converting Rate(TCR) in n-th stage as follows:
$TC_{f}^{n}\;=\;\sum_{y_{f}\in D_{\mathrm{VP}}^{n}}\hskip-6.99997pt\mathbf{1}\bigl(\hat{y}_{f}^{n}=y_{f}\bigr),\quad TCR_{f}^{n}\;=\;\frac{\bigl|TC_{f}^{n}\bigr|}{\bigl|D_{\mathrm{VP}}^{n}\bigr|}\,$
In the above equations, $f\in\{u,v\}$ denotes the sample type (verifiable or unverifiable), and $n\in\{1,2\}$ indicates the stimulation stage. $D_{\text{VP}}^{n}$ represents the input of Vague Perception samples in the $n$-th stage. First stage input $D_{\mathrm{VP}}^{1}$ was extracted from the reasoning dataset. For the second stage, we collect the False Converting part samples as the input, from which we can obtain: $D_{\mathrm{VP}}^{2}=FC^{1}$.
It should be noted that both $TCR_{v}$ and $TCR_{u}$ quantities represent the sum of results from stage 1 and 2, and $TCR^{1}$ denotes the converting rate of both forms data in first stage. To be specific, $TCR_{v}$ = $TCR_{v}^{1}$ + $TCR_{v}^{2}$, and $TCR^{1}$ = $TCR_{v}^{1}$ + $TCR_{u}^{1}$. To integrate performance across both stages, we define Overall Converting Rate (OCR) as follows:
$OCR=\frac{\lvert TC^{1}\cup TC^{2}\rvert}{\lvert D_{\mathrm{vp}}\rvert},\quad TC_{1}\cap TC_{2}=\emptyset$
OCR refers to the portion of under-stimulated samples that turn out to be predicted correctly. By doing this, we can obtain the latent accuracy of LLMs.
2.2 Remind-then-Guide (RtG) Conformity
Label Conformity
We test whether Vague Perception samples display better or worse stability—i.e. whether they are more easily misguided—relative to ordinary settings indicated in Weng et al. (2025). According to previous work, all models demonstrate Conformity. In this part, we first inform the model that it has misunderstood the sample, and then provide a randomly assigned label with varying probabilities, to test whether this misguiding affects its final decision.
We evaluate this phenomenon by setting up one metric for two types of samples: f can be v or u, and we split the results by whether the sample is verifiable or unverifiable. We compute the difference in accuracy between completely wrong and completely correct label guidance and then average the results. Detailed Remind-then-Guide conformity tests on assigning misguiding labels at four given possibilities (100%, 50%, 33.3%, 0%) is available in the Table 891011.
Figure 3: Overall benchmark distribution, which contains 2,840 samples grouped into two forms: story-based (56.3%) and fact-based (43.7%). A detailed breakdown of all topics and “Else” category is provided in the Appendix B.
Reasoning Process (RP) Conformity
We examine how effectively the reasoning chain generated in Stage 1 serves as inspiration or guidance for Stage 2 when it is supplied as additional input. To ensure that the model demonstrates genuine understanding rather than simply being prompted to output the correct answer by accident, we propose two metrics to dive deeper into this phenomenon.
Correct Guiding Rate (CGR) CGR assesses whether the reasoning process output in Stage 1 can assist Stage 2 in stimulating samples to better avoid prior mistakes. Similarly. We remind the LLMs that these explanations may contains flaws because the samples come from the False Converting set. CGR is calculated by the changes of $TCR^{2}$ with and without Reasoning Process.
Reasoning Process Coherence (RPC) This involves re-running Stage 1 with its own prior reasoning process to assess stability and coherence. Similarly, RPC is calculated by changes of $TCR^{1}$ with and without Reasoning Process.
For both metrics, RPC and CGR, below zero indicates that the previous reasoning process is misguiding and results in declined accuracy, while above zero suggests that it is helpful and can guide the LLM towards correct reasoning.
Table 2: Basic info of each dataset.
Dataset Samples Form Writing Source
FLD 600 Conclusion Abstract Gen
FOLIO 640 Conclusion Plain Real
Science 800 Q&A Plain Real
Arts 800 Q&A Plain Real
Table 3: Comparison of six LLMs across four datasets (FLD, FOLIO, Science, Arts), grouped by evaluation settings. RP refers to Reasoning Process under Remind-then-Guide setting. The highest value is marked darker per column.
Model Dataset Vanilla RtG (Label) RtG (RP)
TCR1 TCR2 OCR
Confv
Confu
CGR
RPC
GPT-3.5 FLD 12.61 14.43 25.22 64.25 3.33 -12.30 10.84
FOLIO 14.41 42.10 50.44 16.81 31.97 -42.10 -6.30
Science 21.05 3.33 21.63 19.22 1.96 -3.33 0.59
Arts 26.04 4.00 27.10 14.24 -3.88 0.00 3.12
GPT-4o FLD 18.88 28.19 41.65 8.17 37.38 -23.50 -6.48
FOLIO 37.34 44.40 65.16 9.63 51.90 -44.77 0.00
Science 24.73 7.76 28.96 11.74 4.58 -7.16 -3.68
Arts 18.60 8.24 24.17 8.07 23.31 -8.24 1.14
LLaMA-3.1-8B FLD 41.97 26.03 48.90 50.04 11.04 -18.56 1.09
FOLIO 45.86 11.40 50.84 38.73 13.61 -3.07 5.52
Science 45.94 5.95 50.61 38.60 51.44 -5.95 12.91
Arts 24.26 8.57 25.75 30.99 9.02 -4.77 11.30
LLaMA-3.1-405B FLD 20.65 53.73 41.50 9.35 57.63 -48.28 -17.66
FOLIO 28.76 29.92 49.22 6.11 23.01 -17.69 0.00
Science 32.85 2.50 33.03 9.02 18.50 0.19 0.71
Arts 4.40 0.00 4.40 0.81 28.91 0.00 -0.77
Qwen-2.5-7B FLD 48.31 40.00 68.53 22.06 42.34 -31.67 -5.61
FOLIO 19.73 36.84 47.36 4.92 38.71 -36.84 14.48
Science 15.83 10.00 24.08 55.21 0.00 -2.98 14.59
Arts 29.34 25.37 36.44 43.39 7.20 -13.27 8.47
Qwen-2.5-72B FLD 48.20 9.80 51.21 33.08 9.14 -9.80 1.21
FOLIO 31.51 11.37 38.36 10.65 2.20 -6.82 -1.37
Science 10.81 3.12 11.20 33.42 22.07 0.00 -0.38
Arts 25.00 9.76 25.56 12.52 24.12 0.00 -1.95
3 Experiments
3.1 Benchmarks
To explore the effects of the Vague Perception phenomenon on different types of logical datasets, we classify them based on their labels: (1) Verifiable Data, where the samples have a clear True/False/Unknown label indicating the validity of the reasoning; and (2) Unverifiable Data, where the label is Unknown and samples are unsolvable. Furthermore, as shown in Figure 3, to better understand how LLMs reason across different logical forms, we categorize benchmarks as follows:
Fact-based Benchmarks.
The information is given in atomic relationships, split by unrelated sentences.: FLD (Morishita et al., 2024) supplies three kinds of sample labels annotated as True, False, or Unknown. We randomly sample 600 items: 300 verifiable (True/False) and 300 marked Unknown due to insufficient evidence. FOLIO (Han et al., 2022) is a factual benchmark with the same 3-way labeling. Similarly, we draw 640 samples—320 verifiable samples with T/F labels and 320 Unknown.
Story-based Benchmarks.
Our experimental design considers the differing characteristics of the two forms above. For Fact-based ones, which models tend to process the entire context holistically, so providing additional stimulation can enhance performance. In contrast, we extract samples from ScienceQA (Lu et al., 2022) for story-based tasks—often rooted in professional knowledge domains and relying on models’ inherent knowledge—we minimize guidance to avoid interfering with the model’s reasoning process.
3.2 Experimental Setup
Our evaluation covers both closed-source models: ChatGPT-3.5-turbo and GPT-4o (OpenAI, 2023), and open-source models: LLaMA-3.1 series (Touvron et al., 2023), and Qwen-2.5 series (Bai et al., 2023). Two stimulating strategies are implemented:
$\triangleright$ Concise Stimulation: Basic error notification without correction guidance.
$\triangleright$ Detailed Stimulation: Enhanced detailed approach featuring:(1) Listed Conditions Guidance.(2) Structured Reasoning Constraints. (3) Conditional Verification Steps.The experimental protocol differs between approaches. Concise Stimulation uses a similar format for two-stage error notification process, while Detailed Stimulation employs a different two-stage workflow: a standard prediction followed by an interrogative refinement. The latter stimulates analytical reasoning through questioning and step-by-step error analysis, creating a clear methodological contrast.
To achieve a balance between randomness and coherence, we set all temperature to 0.5, except for 1.0 in ablation study where we want to further examine the quantitative influence between prompt designing and models setup.
In our result detection procedure, the output is first examined for specific words. When such methods are either absent or multiple candidates are detected, GPT-4o is employed to adjudicate and assign the appropriate label. With the exception of the Remind-then-Guide (RtG) setting in the reasoning process, most configurations yield a single label accompanied by few explanations, which facilitates reliable identification.
3.3 Results on Benchmark Converting Rate
As shown in Table 3, we report the overall benchmark Converting Rate, which means three metrics in three systems: True Converting Rate (TCR) for both stages, and Overall Converting Rate (OCR) taking the portion of Vague Perception samples into consideration. The LLaMA and Qwen series show superior Overall Converting Rate (OCR) on small-parameter series, whereas the GPT series shows the opposite trend. We infer that, in general, higher parameter counts correlate with lower OCR. This situation may change due to the distinct yet controversial architecture of GPT-4o.
Table 4: Different Stimulating Method Influence on Converting Rate, where $\mathrm{TCR}^{1}=\mathrm{TCR}_{u}^{1}+\mathrm{TCR}_{v}^{1}$ as we split converting rate by verifiable and unverifiable samples. CS, DS refer to Concise and Detailed Stimulation. Rows shaded gray denote ones used in our framework.
Metrics(%) TCR1 $_{v}^{1} TCR$$_{u}^{1}$ Reliability
FLD (CS) 37.27 10.65 26.62 $\downarrow$
FLD (DS) 18.88 13.64 5.24 $\uparrow$
Arts (CS) 18.60 16.06 2.54 $\uparrow$
Arts (DS) 22.54 14.65 7.89 $\downarrow$
3.4 Results on RtG Conformity Stimulating
As illustrated in the Figure 2, RtG module contains two parts: Label and Reasoning Process (RP) Conformity: We remind the model incorrectly predicted the sample as Unknown before; we then we apply conformity test to verify whether models can inference independently.
Label Conformity
Judging from the table, we observe that even when LLMs are reminded of their previous incorrect answers, they still demonstrate a strong tendency toward conformity. Moreover, distinct conformity patterns emerge between verifiable and unverifiable samples. One key influencing factor appears to be the model size: models with smaller parameter sizes, such as GPT-3.5 and LLaMA-3.1-8B-Instruct, are more inclined to being misguided by verifiable samples with T/F labels reasoning. In contrast, models with larger parameter sizes, such as GPT-4o and LLaMA-3.1-405B, tend to be more easily misguided regarding unverifiable samples with Unknown label. Meanwhile, Qwen series shows distinct yet related features compared with the above, we infer it is because parameters gap between Qwen 7B and 72B is much smaller than the former two series.
Reasoning Process (RP) Conformity
We input the previous reasoning process and explicitly ask the LLM to reflect on its mistakes, the resulting CGR values are mostly below zero, which indicates that even after being reminded of a past error and provided with its own flawed reasoning, the LLM still tends to repeat the same incorrect conclusion, leading to an even larger decline in accuracy.
Our results on RPC reveal that the reasoning process is often highly unstable. Changes in accuracy can largely be attributed to divergence in how the problem is interpreted. Notably, for unverifiable samples, the reasoning process tends to be unreliable — feeding it back into the model typically leads to a further drop in accuracy. This exposes a hidden issue: when LLMs automatically generate solutions for a dataset using a reasoning process, unverifiable problems are particularly susceptible to unreliability. Even for verifiable samples, the reasoning may remain unstable. As a result, the accuracy and reliability of AI-annotated problem-solving methods for datasets remain a problem.
4 Detailed Analysis
4.1 True Converting Rate (TCR) Analysis
As shown in the line chart of Figure 4, by comparing TCR1 and TCR2 across different benchmarks, we observe clear divergences: This indicates that many samples mis-predicted in Stage 1 can be self-corrected in Stage 2, where the model refines its initial output. In contrast, Story-based TCR1 shows a substantially lower TCR2. Whereas LLaMA-3.1-8B demonstrates much higher TCR1.
Thus, distinct focus strategies are recommended for multi-stage frameworks for the following LLMs: (i) For GPT-4o, when facing Fact-based benchmarks, broader exploratory freedom is acceptable in Stage 1. “making mistakes at first is acceptable", since later refinement proves beneficial; For Story-based benchmarks, the improvement of the subsequent stage is limited, so initial accuracy should be emphasized (ii) For LLaMA-3.1-8B, it is always better to “do it right the first time" .
4.2 Benchmark-wise Vague Perception
As each model differs in its reasoning ability, we first quantify the percentage of data in which models induce Vague Perception. The Bar chart in blue Fig 4 reports the percentage for GPT-4o—the most widely used LLM at present—while complete results for the remaining models in Appendix 7.
We can conclude that the Vague Perception phenomenon is prevalent across all benchmarks. For GPT-4o, FOLIO exhibits a relatively lower percentage. For FLD and Arts—which focus purely on language understanding, the model demonstrates greater uncertainty regarding explaining why certain samples are labeled as Unknown; In stark contrast, Science—which comprises Physics and Biology topics—faces significant challenges even when the samples are verifiable; For LLaMA-3.1-8B, there is a significant rise in the percentage of verifiable samples predicted as Unknown, but the ability to comprehend unverifiable samples is stronger.
[…]
4.3 Stimulating Style Comparison
[…]
Furthermore, we observe a latent trade-off: as the ability to resolve verifiable samples declines, the model becomes more confident on unverifiable ones. We attribute this to a shift in the model’s reasoning habits.
4.4 Ablation Study
From the Stage 1 output, we isolated the Unexcited samples—those that failed to even generate a complete logic chain—and fed them back through the same Stage 1 prompting protocol. Both LLaMA-3.1-8b-Instruct and GPT-4o are set with 1.0 temperature to ensure randomness.
As Figure 4 shows, this second pass produces virtually no additional conversions: more than 95% of the samples remain unchanged. This finding demonstrates that Vague Perception is largely immune to repeated, identical n-shot prompting; applying the same stimulation twice is therefore futile. Consequently, any effective correction must rely on a distinct stage 2 rather than reiterating stage 1, adjusting temperature alone without revising stimulation style can be limited.
Figure 5: Distribution of the root causes behind GPT-4o’s Vague Perception errors across all benchmarks (left). Qwen-2.5-7B and 72B model False and Unexcited Converting Rate comparison (Right). “FU”: Fact Understanding, “RG”: Reasoning Gap, and “Else” aggregates residual minor factors.
4.5 Baseline Comparison
In our work, we primarily analyze the “waken ability” with respect to Vague Perception samples by computing the proportion of truly converted samples in both stages ($TC^{1}$ and $TC^{2}$) relative to the percentage of Vague Perception samples; however, this constitutes only one component, we combine this component with the vanilla accuracy by adding them together, thus we gain an empirical bound of reasoning accuracy—namely, the maximum reasoning capability that LLMs can achieve without additional training or data-processing. The empirical accuracy of WakenLLM is defined by two parts: (1) the proportion of samples correctly reasoned by the LLM, and (2) the truly converted samples across both stages ($TC^{1}$, $TC^{2}$) that are potentially solvable. Acc(Vanilla) denotes the reasoning accuracy under direct prompt setting, without any intermediate interpretation. Here, D represents the complete dataset, whereas $D_{VP}$ only represents a shortlisted ones that induces Vague Perception.
Acc(WakenLLM) =
$\underbrace{\text{Acc(Vanilla)}}_{\text{Explicit Ability}}\;+\;\underbrace{|TC^{1}\cup TC^{2}|/|D|}_{\text{Latent Ability}}$
As illustrated in the left section of the line chart in Figure 5, most baseline methods exhibit a substantial reasoning gap compared to our pipeline, indicating most current frameworks failed to trigger models’ full potential. Refer to the Table 6 for detailed experimental results.
4.6 Phenomenon Exploration
To probe the root causes of GPT-4o’s mistakes, we annotate each error sample with one of three labels:
$\triangleright$ Fact Understanding (FU) – The model does not correctly retrieve, align, or synthesize the relevant facts stated in the passage.
$\triangleright$ Reasoning Gap (RG) – The model retrieves the facts but fails to perform the intermediate logical operations that connect facts to the hypothesis.
$\triangleright$ Else – For the two above generally covers 80% or more scenarios. And there is a catch-all category for other less frequent factors. Notably, issues related to Excessive Caution are particularly prominent in this part.
Excessive Caution
Our experiments reveal that Excessive Caution—instances where the model responds with Unknown despite sufficient evidence—is nearly absent in Fact-based benchmarks but increases from 2.4% in Science to 5.4% in Arts. We attribute this rise to the narrative complexity of Story-based tasks: unlike the atomic and structured format of Fact-based benchmark, passages in Arts often embed literary allusions and background knowledge, heightening epistemic uncertainty.
Table 5: Performance comparison of Degradation Analysis Phenomenon between GPT‑4o and GPT‑3.5.
Models GPT‑4o GPT‑3.5
Metrics(%) TCR2 Deg TCR2 Deg
[…]
4.7 Degradation Analysis
Within the pipeline, we don’t exclude the possibility that previous incorrectly predicted samples, FC2 from Stage 1 may be converted to Unknown, which should be denoted as UC2. As it is reasonable that at first, the model may generate an incorrect answer, but after further stimulation, due to its limited ability, it fails to arrive at the correct answer and can only output Unknown. We elaborate this as:
$Deg=\frac{\left|FC^{1}\setminus\left(FC^{2}\cup TC^{2}\right)\right|}{\left|FC^{1}\right|},\ FC^{2}\cap TC^{2}=\emptyset$
As shown in Table 5, only a small proportion of samples exhibit this phenomenon. However, in the case of GPT-4o on the reasoning-based Science dataset, the issue becomes more serious.
As for the Science dataset, GPT-4o achieves the highest TCR2 among all four datasets. Coupled with higher Deg value, this suggests that when confronted with problems it previously answered incorrectly—especially those beyond its capability—GPT-4o tends to respond with an Unknown label. In contrast, GPT-3.5-turbo is much more likely to provide a definitive but incorrect answer instead of admitting uncertainty.
This can be further proved by the uniform distribution of UCR and FCR in Appendix D acquired similar to formula TCR, denoting Unknown and False Converted rate in the Vague Perception input. We compare answering of Qwen-2.5 series (7B and 72B model), as shown in the radar-chart in Figure 5: (1) Models show distinctive yet uniform preferences among four different forms of dataset. For FLD and FOLIO, models are much more willing to output a false answer rather than admitting Unknown, this is even more serious for FOLIO dataset, where model likes to “show off" when facing easier ones, as FOLIO achieves the highest accuracy among above four; Story-based dataset, like Science and Arts are exactly the opposite. (2) The higher the parameters, the more likely models are to output Unknown when inputting problems beyond LLMs’ ability.
5 Conclusion
We present WakenLLM, a fine-grained framework for evaluating the Vague Perception phenomenon of LLMs. The framework supplies targeted cues that stimulate models to reflect on reasoning, substantially raising the empirical accuracy. Experiments show that while LLMs give up and answer Unknown on the first try, well-designed stimulation reliably steers them to the correct label and uncovers a persistent tendency toward conformity even after explicit reminding. These insights suggest training and evaluation protocols that value a model’s capacity for potential instead of the one-shot performance alone. Even though significant improvements in inner structures and parameter optimizations of LLMs may be limited in the near future, however, our findings suggest that current methods are still far from triggering LLMs’ full potential, and addressing this problem can bring unique prospective. Our research offers high underlying benefits regarding models’ potential.
6 Limitations
Even though our work has helped expose the empirical reasoning potential of LLMs, however, our final metrics are all calculated regarding labels, and outputting true labels with flawed reasoning process can be a common problem. How to fundamentally quantify their true ability with true reasoning process remains a problem in the future.
7 Ethical Considerations
Our research utilizes publicly available datasets (FLD, FOLIO, ScienceQA) containing synthetic logical reasoning problems without personal information. We acknowledge potential biases in our SFT and DPO training processes and implement atomic reasoning criteria to ensure fair evaluation. All training scripts, evaluation code, and prompt templates are made publicly available for reproducibility. We use LoRA fine-tuning to reduce computational requirements and focus on improving logical reasoning quality rather than enabling harmful applications. Our work adheres to academic integrity standards and contributes to responsible AI development through more interpretable reasoning.
Appendix A Baseline Details
Chain-of-Thought (CoT)
The CoT prompting method guides large language models (LLMs) to generate intermediate reasoning steps before producing a final answer. It improves multi-step reasoning performance by making the inference process explicit. This technique is particularly effective in arithmetic and logic-heavy tasks, as shown in prior works such as Wei et al. (2022b).
Few-shot
Few-shot prompting involves presenting the model with a few annotated examples in the input to demonstrate the task format. It enables in-context learning without any gradient updates. Despite its simplicity, few-shot prompting serves as a strong baseline in many LLM applications (Brown et al., 2020).
Tree-of-Thought (ToT)
Tree-of-Thought extends the CoT framework by exploring multiple reasoning paths in a tree structure and evaluating intermediate steps. It mimics a planning process that considers alternative possibilities, significantly enhancing reasoning robustness and allowing backtracking or reevaluation of incorrect paths (Yao et al., 2023a).
Plan & Solve
Plan-and-Solve separates the reasoning process into two stages: first planning a high-level outline or decomposition of the task, then solving sub-tasks based on that plan. This two-stage strategy enhances controllability and consistency in reasoning, and has shown strong performance across complex benchmarks (Wang et al., 2023a).
Supervised Fine-Tuning (SFT)
Supervised Fine-Tuning (SFT) is the standard method for aligning LLMs by training on curated human-annotated question-answer pairs. Unlike prompt-only methods, SFT optimizes the model weights directly through gradient updates to produce desirable outputs. It forms the foundation of instruction-tuned models like InstructGPT and serves as a widely-used supervised baseline for alignment tasks (Ouyang et al., 2022). The common consensus is that SFT provides the essential starting point for alignment, but by itself is insufficient to fully capture complex human preferences, often requiring further preference optimization or reinforcement learning. The optimized formula is below:
$\mathcal{L}_{\mathrm{SFT}}(\theta)=-\sum_{i=1}^{N}\log\pi_{\theta}(y_{i}\mid x_{i}),$
Direct Preference Optimization (DPO)
Direct Preference Optimization (DPO) is a reinforcement learning-free approach that fine-tunes LLMs using preference data. Given pairs of outputs (e.g., preferred vs. dispreferred), DPO directly optimizes the log-likelihood ratio between them, providing alignment benefits similar to reinforcement learning methods like PPO while avoiding their instability and complexity (Rafailov et al., 2023). The common consensus is that DPO offers a simpler and more stable alternative to reinforcement learning-based methods, making it a practical choice for preference alignment, though it may still struggle with capturing nuanced or long-term user preferences. The formula can be concluded as below:
$\begin{split}\mathcal{L}_{\mathrm{DPO}}(\theta)=-\mathbb{E}_{(x,y^{+},y^{-})}\Bigl[\log\sigma\!\Bigl(\beta\bigl(\log\tfrac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\mathrm{ref}}(y^{+}\mid x)}\\ -\log\tfrac{\pi_{\theta}(y^{-}\mid x)}{\pi_{\mathrm{ref}}(y^{-}\mid x)}\bigr)\Bigr)\Bigr]\end{split}$
A.1 Experiment Setup
We choose four LLMs: GPT-4o, GPT-3.5-turbo-1105, LLaMA-3.1-8B and Qwen-2.5-7B for our detailed baseline experiments. We categorize our baseline methods into two kinds: training and non-training, in order to see whether the Vague Perception phenomenon can be solved with or without parameter revision. For model training and open-source model deployment, the experiments ran on four Tesla V100-SXM2-32GB GPUs.
A.2 Training Methods Details
As explained before in Appendix A.1, we elaborate our facilities and setup for training. In this section, we demonstrate our details for the training.
For SFT (Supervised Fine-Tuning), we extract 1,000 training samples from exactly same dataset, the prompt instruction can be found in Appendix D.2.
For DPO (Direct Preference Optimization), we apply LLM Judge to help select the reasoning process with many atomic steps as possible. This includes two parts: (1) Prompt Template that includes task guiding details (e.g., definition of atom steps) and task instructions (2) Testing cases that can perfectly prove and verify which response is better given the metrics and requirements. All information above can be found in Appendix D.2.
Appendix B Benchmark Details
FLD
Following Morishita et al. (2024), FLD is a fully synthetic benchmark for multi‑step logical inference. Formal proof chains of depth 0–15 are automatically generated and then verbalized into natural‑language premises plus a hypothesis. Each instance is labeled Proved, Disproved, or Unknown, making the dataset ideal for assessing whether a model can (i) draw correct conclusions and (ii) abstain when evidence is insufficient.
FOLIO
FOLIO (Han et al., 2022) contains 1,430 human‑written first‑order‑logic stories that bridge everyday commonsense with formal theorem proving. Every narrative provides its gold logical forms and a hypothesis whose truth value (True, False, or Unknown) is automatically certified. The official split is 1 001/226/203 for train/dev/test, offering fewer but richer problems than FLD and requiring explicit refusals when the passage leaves the hypothesis undecided.
ScienceQA
ScienceQA (Lu et al., 2022) is a multi-modal multiple‑choice benchmark with 21,208 elementary‑ and middle‑school questions across 26 topics. To create an abstention setting comparable to the logic corpora, we select questions from Physics and Biology (science) as well as Writing‑Strategies and Figurative‑Language (arts). For 50% of the chosen items we delete two random supporting sentences with GPT‑4o, then manually verify that no unique answer remains; these items are re-labeled UNKNOWN. The final corpus therefore mixes standard A–D supervision with an explicit refusal class.
Topic Analysis
Our comprehensive dataset consists of four distinct collections, each representing different domains of logical reasoning. The FLD (First-Order Logic Dataset) contains 600 samples across 11 topics, with Philosophy being the dominant subject (545 samples, 90.8%), followed by Logic (37 samples, 6.2%), and 9 other topics with minimal representation. The FOLIO dataset comprises 640 samples distributed across 7 topics, including Arts (146 samples, 22.8%), Science & Technology (130 samples, 20.3%), Sociology (114 samples, 17.8%), Logic (102 samples, 15.9%), Sports (85 samples, 13.3%), Geography (39 samples, 6.1%), and Architecture (24 samples, 3.8%). The ScienceQA dataset contains 800 samples evenly split between Biology and Physics (400 samples each, 50%). Finally, the Arts dataset includes 800 samples with Literature as the primary topic (526 samples, 65.8%), followed by Education (190 samples, 23.8%), Philosophy (77 samples, 9.6%), Linguistics (4 samples, 0.5%), and Psychology (3 samples, 0.4%).
The dataset distribution reveals a hierarchical structure: at the top level, we distinguish between Fact-based reasoning (43.66%) and Story-based reasoning (56.34%). The four major categories are distributed as follows: FLD (21.13%), FOLIO (22.53%), Science (28.17%), and Arts (28.17%). This balanced distribution ensures comprehensive coverage of different reasoning paradigms, from formal logical reasoning in FLD to narrative-based reasoning in Arts, providing a robust foundation for training and evaluating logical reasoning capabilities across diverse domains.
Unknown Labeling
For the first FLD and FOLIO, the Unknown label is originally included in their annotations. For ScienceQA, we randomly select half of the samples and employ GPT-4o to randomly delete two sentences from the given facts. We then re-label these modified samples as Unknown. Additionally, we perform manual verification to ensure that each relabeled sample is semantically consistent with its assigned label.
Ethical Statement
All datasets are distributed for non‑commercial research under permissive licenses (FLD: MIT; FOLIO: research‑only; ScienceQA: CC‑BY‑NC‑SA 4.0). Our sentence‑removal augmentation introduces no personal data and was human‑verified to avoid unsafe or misleading content.
Figure 6: Scope of Vague Perception. This phenomenon underlies a wide range of reasoning failures. By identifying and sub-categorizing cases that are potentially solvable, it reveals a substantial portion of failures that stem from Vague Perception rather than fundamental incapacity. Addressing this issue can significantly advance fine-grained reasoning and improve the overall trustworthiness of LLMs.
Model Methods Accuracy (%)
FLD FOLIO Science Arts
GPT-3.5 Direct 50.17 40.63 44.63 29.25
CoT 49.67 42.34 44.13 31.00
Few-shot 49.67 37.50 43.75 33.88
Tree-of-Thought 33.50 43.91 36.00 28.25
Plan & Solve 42.83 45.47 33.63 26.38
WakenLLM 60.26 46.68 48.38 35.04
GPT-4o Direct 51.00 69.53 43.00 33.88
CoT 54.10 69.53 43.12 31.87
Few-shot 58.30 74.40 47.62 50.50
Tree-of-Thought 52.83 70.62 43.38 29.25
Plan & Solve 52.17 69.84 42.38 30.25
WakenLLM 71.79 75.74 56.25 44.64
LLaMA-3.1-8B Direct 39.17 51.56 49.50 47.50
CoT 40.67 51.88 46.50 43.25
Few-shot 33.33 51.25 48.25 39.13
Tree-of-Thought 38.33 47.66 40.63 30.00
Plan & Solve 35.50 52.81 38.75 26.75
SFT 49.17 50.47 46.50 49.12
DPO 50.33 48.60 48.50 49.62
WakenLLM 61.50 65.47 62.72 53.43
Qwen-2.5-7B Direct 41.83 57.81 44.12 47.12
CoT 43.83 60.62 42.75 46.62
Few-shot 44.50 58.75 38.00 42.25
Tree-of-Thought 39.83 57.19 37.88 45.88
Plan & Solve 40.67 56.88 37.62 39.62
SFT 48.67 57.34 49.62 43.12
DPO 49.33 54.21 48.38 44.13
WakenLLM 48.06 65.55 47.56 54.37
Table 6: Accuracy comparison of multiple baselines for four LLMs evaluated on four datasets. Each block represents a distinct model series. The highest accuracy is marked black.
Model FLD FOLIO Science Arts
V U V U V U V U
GPT-3.5 72.00 8.00 10.00 24.69 40.00 2.75 22.50 1.50
GPT-4o 11.00 88.00 7.81 11.25 68.75 22.75 28.25 60.75
LLaMA-3.1 (8B) 76.33 15.00 54.69 1.88 49.25 3.00 46.00 4.50
LLaMA-3.1 (405B) 22.67 100.00 20.94 99.69 50.00 89.25 9.00 99.00
Qwen-2.5 (7B) 20.67 9.00 10.00 13.75 60.00 0.00 52.50 6.50
Qwen-2.5 (72B) 49.67 5.67 15.31 7.50 54.75 10.00 27.75 36.25
Table 7: The percentage of LLMs demonstrating the Vague Perception phenomenon among six LLMs. Split by two different root causes: (1) V: Verifiable samples with T/F labels, but the model output Unknown. (2) U: Unverifiable Samples with Unknown label, but the model output Unknown simply because of incapacity rather than fully understand this is objectively unsolvable.
Search-Based Planning Frameworks
Instead of generating a single chain, Tree-of-Thought (ToT) views partial solutions as nodes in a tree and performs breadth-first search with value heuristics (Yao et al., 2023b). Decomposed Prompting turns a complex task into callable sub-prompts that can be solved (and verified) independently, achieving strong zero-shot gains on compositional reasoning benchmarks (Khot et al., 2023). Reasoning-via-Planning (RAP) couples an LLM world-model with Monte-Carlo Tree Search for long-horizon tasks (Hao et al., 2023). Cochain (Chain-of-Collaboration Prompting) extends this line by orchestrating multiple lightweight agents whose specialized sub-chains are merged via a knowledge-graph plan, outperforming single-agent CoT on workflow-style tasks while using fewer tokens (Zhao et al., 2025).
Appendix C Related Work Details
Chain-of-Thought (CoT) prompting (Wei et al., 2022a) first showed that furnishing step-wise exemplars markedly improves LLM performance on arithmetic and commonsense problems. Follow-up work refines this paradigm: scratchpads (Nye et al., 2021), least-to-most prompting (Zhou et al., 2022), and self-consistency (Wang et al., 2023c). More recently, multi-agent and symbolic reasoning approaches like multiagent debate (Du et al., 2024), symbolic chain-of-thought (Xu et al., 2024), and neurosymbolic reasoning with logic provers (Olausson et al., 2023) further strengthen factuality and logical consistency.
Tool-Augmented Reasoning
Program-Aided LMs (Gao et al., 2022) delegate arithmetic or logical computation to Python. ReAct (Yao et al., 2022) interleaves reasoning and external tools. CRITIC (Gou et al., 2024) validates intermediate answers with execution tools. Chain-of-Verification (Dhuliawala et al., 2024) reduces hallucinations by verifying sub-claims against evidence.
Iterative Self-Refinement and Critique
Self-Refine (Madaan et al., 2023) establishes a generate–criticise–rewrite loop. Ambiguity handling and uncertainty decomposition are also explored: AmbigQA (Min et al., 2020), Decomposing Uncertainty (Hou et al., 2024), and rank-based calibration (Huang et al., 2024).
Datasets for Reasoning and Verification
Classic benchmarks like BoolQ (Clark et al., 2019), SQuAD 2.0 (Rajpurkar et al., 2018), FEVER (Thorne et al., 2018), and HotpotQA (Yang et al., 2018) stress-test fact verification, ambiguity, and multi-hop reasoning.
Active Prompting
Active-Prompt (Diao et al., 2024) selects informative exemplars to boost performance under budget constraints. Meanwhile, Ling et al. (2025) denotes different prompt coverage leads to various effects.
Faithfulness and Safety
Beyond correctness, methods ensure interpretability and reliability, such as Faithful-CoT (Lyu et al., 2023), symbolic verification (Xu et al., 2024), and multi-agent debate (Du et al., 2024).
Appendix D Formula Details
D.1 False and Unknown Converting
As demonstrated in 2.1, we denote the notion of TCR. Likewise, we define the concepts regarding false converting and unknown converting similarly:
$FC_{f}^{n}\;=\;\sum_{y_{f}\in D_{\mathrm{VP}}^{n}}\hskip-6.99997pt\mathbf{1}\bigl(\hat{y}_{f}^{n}\neq y_{f},\;\hat{y}_{f}\notin U\bigr)$
$FCR_{f}^{n}\;=\;\frac{\bigl|FC_{f}^{n}\bigr|}{\bigl|D_{\mathrm{VP}}^{n}\bigr|}$
$UC_{f}^{n}\;=\;\sum_{y_{f}\in D_{\mathrm{VP}}^{n}}\hskip-6.99997pt\mathbf{1}\bigl(\hat{y}_{f}^{n}\in U,\;y_{f}\notin U\bigr)$
$UCR_{f}^{n}\;=\;\frac{\bigl|UC_{f}^{n}\bigr|}{\bigl|D_{\mathrm{VP}}^{n}\bigr|}$
Note that, Unknown Converting (UC) is still categorized into Vague Perception phenomenon: Outputting unknown when facing verifiable samples. We can tell the following principles for any stage, n belongs to 1 or 2:
$UCR_{f}^{n}+FCR_{f}^{n}+TCR_{f}^{n}\;=\;100\%$
$UC_{f}^{n}\cup FC_{f}^{n}\cup TC_{f}^{n}\;=\;D_{VP}^{n}$
$UC_{f}^{n}\cap FC_{f}^{n}\cap TC_{f}^{n}\;=\;\emptyset$
Our metrics systematically define and categorize the flowing of output distribution by means of above metrics.
D.2 Overall Converting & Latent Ability
We clarify the relationship between Overall Converting Rate (OCR) in Section 2.1 with the latent ability in 4.5, as we depicted, by comparing True Converting samples (TC) with $D_{VP}$, we can get a empirical ability of how many samples can be eventually converted to right labels within all Vague Perception samples, whereas the latter indicates overall accuracy gain among the all samples D in a broader aspect. By choosing different objects to compare with, we can gain insights not only in the phenomenon alone, but also the general LLMs’ reasoning potential.
Table 8: Detailed Remind-then-Guide (RtG) setting of GPT-3.5 and GPT-4o. M denotes Misguiding Rate: the percentage of input labels are replaced with wrong one. We break down $TCR^{1}$ into verifiable and unverifiable data: $TCR_{v}^{1}$ and $TCR_{u}^{1}$
Model Dataset M (%) TCR1 (%) TCR2 (%)
V U V U
GPT-3.5 FLD 100 5.83 3.33 1.67 0.00
66.67 20.83 6.67 16.67 2.78
50 27.50 6.00 40.54 10.81
0 54.17 10.00 81.82 0.00
FOLIO 100 2.70 0.90 0.00 0.00
66.67 6.31 10.81 3.80 12.66
50 9.91 22.52 3.23 20.97
0 18.92 38.74 17.39 26.09
Science 100 19.88 0.00 3.57 7.00
66.67 31.58 1.17 3.70 7.00
50 30.41 1.75 7.02 11.00
0 43.86 2.92 18.03 8.00
Arts 100 22.92 0.00 0.00 15.00
66.67 21.88 1.04 0.00 6.00
50 34.38 3.00 9.38 6.00
0 37.50 4.17 13.89 3.00
GPT-4o FLD 100 15.27 10.09 0.39 4.65
66.67 19.60 12.39 13.74 4.09
50 21.04 17.00 2.33 34.88
0 26.22 27.09 5.81 62.40
FOLIO 100 5.85 6.73 0.00 6.06
66.67 11.11 21.05 1.14 25.38
50 13.45 32.75 4.17 37.12
0 20.18 54.09 4.92 62.50
Science 100 8.42 5.79 1.54 25.00
66.67 11.84 7.37 4.62 28.00
50 14.21 7.37 6.15 24.00
0 20.26 8.95 13.08 32.00
Arts 100 11.83 4.51 1.06 24.00
66.67 14.37 7.32 4.58 38.00
50 14.65 10.70 5.99 45.00
0 17.75 15.21 11.27 60.00
Table 9: Detailed Remind-then-Guide (RtG) setting of LLaMA-3.1-8B and LLaMA-3.1-405B. M denotes Misguiding Rate: the percentage of input labels are replaced with wrong one. We break down $TCR^{1}$ into verifiable and unverifiable data: $TCR_{v}^{1}$ and $TCR_{u}^{1}$. "/" denotes there is no input for stage 2, as false converted samples ($FC^{1}$ from stage 1 don’t exist.
Model Dataset M (%) TCR1 (%) TCR2 (%)
V U V U
LLaMA3.1 (8B) FLD 100 0.36 2.92 0.00 2.36
66.67 21.90 6.57 22.58 16.13
50 39.05 11.00 31.25 10.94
0 75.18 16.42 / /
FOLIO 100 2.76 0.55 1.00 0.00
66.67 26.52 0.55 25.45 1.82
50 41.44 0.00 39.13 6.52
0 81.22 2.76 0.00 25.00
Science 100 0.96 1.44 5.10 0.00
66.67 28.23 1.44 22.81 7.00
50 37.32 3.83 45.65 2.00
0 83.25 4.31 0.00 100.00
Arts 100 22.77 1.49 14.61 3.00
66.67 33.17 2.97 29.41 7.00
50 42.57 3.96 34.72 6.00
0 60.89 4.95 38.46 18.00
LLaMA3.1 (405B) FLD 100 1.90 0.82 0.00 0.31
66.67 6.52 16.30 5.22 30.22
50 8.42 28.00 5.41 35.14
0 17.66 54.62 2.94 61.76
FOLIO 100 4.40 18.91 1.00 7.91
66.67 8.29 28.50 2.62 17.03
50 8.55 33.94 2.43 21.36
0 14.51 48.19 2.82 24.65
Science 100 7.72 15.00 2.27 35.00
66.67 13.64 21.54 3.25 37.00
50 16.70 25.31 4.06 42.00
0 25.13 38.78 2.89 48.00
Arts 100 3.24 0.93 0.24 8.00
66.67 3.24 5.09 0.00 13.00
50 4.40 6.71 0.26 15.00
0 5.09 10.42 0.00 56.00
Table 10: Detailed Remind-then-Guide (RtG) setting of Qwen-2.5-7B and Qwen-2.5-72B. M denotes Mis-guiding Rate: the percentage of input labels are replaced with wrong one. We break down $TCR^{1}$ into verifiable and unverifiable data: $TCR_{v}^{1}$ and $TCR_{u}^{1}$
Model Dataset M (%) TCR1 (%) TCR2 (%)
V U V U
Qwen-2.5-7B FLD 100 14.61 1.12 2.08 2.08
66.67 28.09 5.62 9.52 14.29
50 34.83 4.00 17.65 29.41
0 52.81 7.87 8.00 80.00
FOLIO 100 15.79 3.95 2.00 7.84
66.67 19.74 13.16 0.00 33.33
50 25.00 11.84 0.00 23.08
0 27.63 23.68 0.00 65.52
Science 100 16.67 0.00 2.82 0.00
66.67 27.50 0.00 23.73 0.00
50 32.08 0.00 39.34 0.00
0 50.83 0.00 79.07 0.00
Arts 100 16.95 1.27 6.67 4.00
66.67 31.78 4.24 21.43 10.00
50 38.14 3.39 35.00 14.00
0 55.93 6.36 52.46 13.00
Qwen-2.5-72B FLD 100 13.25 1.20 2.99 2.99
66.67 38.55 2.41 20.75 5.66
50 42.17 3.00 19.44 13.89
0 77.71 4.82 5.88 17.65
FOLIO 100 19.18 0.00 2.00 6.98
66.67 24.66 0.00 5.41 8.11
50 30.14 0.00 10.81 5.41
0 42.47 1.37 0.00 10.00
Science 100 8.88 0.00 0.00 10.00
66.67 17.76 0.00 13.79 16.09
50 22.78 0.00 20.69 25.29
0 37.84 0.00 37.88 55.00
Arts 100 13.28 2.34 3.33 3.00
66.67 16.80 3.91 9.88 13.00
50 21.09 4.30 13.10 28.00
0 24.22 9.38 17.42 45.00
Table 11: Six LLMs’ $TCR^{1}$ and $TCR^{2}$ break down into verifiable and unverifiable samples: $TCR_{v}^{1}$, $TCR_{u}^{1}$, $TCR_{v}^{2}$, $TCR_{u}^{2}$, among all four datasets.
Model Dataset TCR1 (%) TCR2 (%)
V U V U
GPT-3.5 FLD 12.61 0 9.28 5.15
FOLIO 14.41 0 4.21 37.89
Science 21.05 0 3.33 0
Arts 22.92 3.12 4.00 0
GPT-4o FLD 13.64 5.24 1.32 26.87
FOLIO 33.73 3.61 0.37 44.03
Science 21.05 3.68 0.78 6.98
Arts 16.06 2.54 0.35 7.89
LLaMA-3.1-8B FLD 33.21 8.76 4.11 21.92
FOLIO 44.20 1.66 10.13 1.27
Science 44.98 0.96 5.95 0
Arts 18.81 5.45 8.57 0
LLaMA-3.1-405B FLD 5.43 15.22 0 53.73
FOLIO 9.59 19.17 0 29.92
Science 10.23 22.62 2.50 0
Arts 4.17 0.23 0 0
Qwen2.5-7B FLD 47.19 1.12 0 40.00
FOLIO 14.47 5.26 1.75 35.09
Science 15.83 0 10.00 0
Arts 23.31 5.93 25.37 0
Qwen2.5-72B FLD 46.39 1.81 0 9.80
FOLIO 30.14 1.37 4.55 6.82
Science 10.81 0 3.12 0
Arts 18.75 6.25 9.76 0
Prompt Template 1: Concise Stimulation Here is the Hypothesis: {} Now These are the facts: {} The answer is {} The data where you previously incorrectly inferred — pay more attention and avoid mistakes. Think twice before outputting, as the answers can be reasoned out because of objective factors. Subjective ability limits should be overcome. These can be reasoned out either whether the answer to the question is __PROVED__, __DISPROVED__, or __UNKNOWN__ — simply because of insufficient information. Output: Conclusion: __PROVED__, __DISPROVED__, or __UNKNOWN__
Prompt Template 2: Detailed Stimulation Here is the Hypothesis: {} Now These are the facts: {} Task: Determine if the hypothesis is logically proved by the facts. You must choose one of the following labels based on logical reasoning: __PROVED__: Facts clearly support the hypothesis; all necessary connections are present, no contradictions, and support is unambiguous. __DISPROVED__: Facts clearly contradict the hypothesis; no supporting facts exist, and the contradiction is logically sound. __UNKNOWN__: Only use when facts are truly insufficient, not due to subjective limitations or unclear reasoning. Critical Rules: Think carefully before making your decision. These cases can be reasoned out objectively. Never output __UNKNOWN__ due to performance limits. Output exactly one of: __PROVED__, __DISPROVED__, or __UNKNOWN__. Do not include any additional text or explanation. Reminder: You should NOT output __UNKNOWN__ due to your own ability limitations. Try harder when the reasoning is difficult. Your output must be exactly one of: __PROVED__  __DISPROVED__  __UNKNOWN__
Prompt Template 3: Output Reasoning Process for Remind-then-Guide (RtG) Setting Here is the Hypothesis: {} Now These are the facts: {} The answer is {} The Data where you previously incorrectly inferred — pay more attention and not to make mistakes. Think twice before outputting, as they can be reasoned out because of objective factors. Subjective ability limits should be conquered. Please follow these steps: 1. First, carefully analyze the facts, the question and the answer 2. Then, explain your reasoning process step by step 3. Finally, conclude with your answer Important guidelines: If the facts logically support the answer to the question, output __PROVED__ If the facts logically contradict the answer to the question, output __DISPROVED__ If the facts are insufficient to prove or disprove, output __UNKNOWN__ Think twice before you output __UNKNOWN__; it is prohibited due to subjective limitations Please format your response as follows: Reasoning Process: [your step-by-step reasoning here] Conclusion: __PROVED__, __DISPROVED__, or __UNKNOWN__
Prompt Template 4: Remind-then-Guide (RtG) Setting – Bringing Reasoning Process into Consideration Here is the Hypothesis: {} Now These are the facts: {} The answer is {} You previously incorrectly inferred the label — it could be worked out by the facts, but you worked out the wrong label. Now your task is to determine if the hypothesis is __PROVED__, __DISPROVED__, or __UNKNOWN__ by the facts. Important guidelines: If the facts logically support the question to the answer, output __PROVED__. If the facts logically contradict the question to the answer, output __DISPROVED__. If the facts are insufficient to prove or disprove the relation, output __UNKNOWN__. Do not output any additional text. Think twice before you output __UNKNOWN__; it is prohibited to output __UNKNOWN__ due to subjective factors like your ability. Previous reasoning process: {}
Prompt Example: Input and output of Detailed Stimulation Facts: A segno is Sardinian. Something that does not lace discredited neither is a Wagner nor remembers Icteria. If something is a Koln then the proposition that it is not a segno and it is not a morphophoneme is incorrect. If that submucosa is a morphophoneme then Joseluis is a segno. That medroxyprogesterone is Sardinian. If that napa is not a Pseudemys then it broadens Aesculapius and it hulkings. That napa is not a Pseudemys. If Joseluis is a morphophoneme that submucosa is a segno. If Joseluis is a segno this is a morphophoneme. This mercury is a morphophoneme. If that napa is classified as a segno that napa is a morphophoneme. Something is a segno if it is Sardinian and not a Koln. A segno is a morphophoneme. If somebody broadens Aesculapius it is Sardinian. That submucosa is a morphophoneme. If something that is not a Wagner does not remember Icteria then that it is not a kind of a Koln is correct. fact17: Joseluis is a morphophoneme if that submucosa is a segno. fact18: Joseluis is a kind of a morphophoneme if that submucosa is Sardinian. Conclusion: Joseluis is not Sardinian. LLM Output: __UNKNOWN__ True Answer: __DISPROVED__
Example: Input and output under Concise stimulation Facts: Organisms, including people, have both inherited and acquired traits. Inherited and acquired traits are gained in different ways. Inherited traits are passed down from biological parents to their offspring through genes. Genes are pieces of hereditary material that contain the instructions that affect inherited traits. Offspring receive their genes, and therefore gain their inherited traits, from their biological parents. Inherited traits do not need to be learned. Acquired traits are gained during a person’s life. Some acquired traits, such as riding a bicycle, are gained by learning. Other acquired traits, such as scars, are caused by the environment. Parents do not pass acquired traits down to their offspring. Question: What information supports the conclusion that Clara acquired this trait? Conclusion: Clara can cook food over a fire. LLM Output: __UNKNOWN__ True Answer: __DISPROVED__
Example: Input and output under Remind-then-Guide (RtG) setting acquiring reasoning process Facts: Every object is made of one or more materials. A material is a type of matter. Wood, glass, metal, and plastic are common materials. A material has different properties. A material’s properties tell you how it looks, feels, tastes, or smells. Question: Which is stickier? Choices: A. syrup B. silk necktie Conclusion: syrup LLM Reasoning Output: The facts define that materials have properties related to how they feel. Stickiness is a tactile property. Syrup is widely known to be sticky, while a silk necktie is smooth and not sticky. Thus, based on general knowledge of material properties, syrup is stickier than a silk necktie. LLM Output: __PROVED__ True Answer: __PROVED__
SFT Instruction Format SFT Data Format: instruction: "You are a helpful assistant. You are given a conclusion and a set of facts. You need to determine if the conclusion is __PROVED__, __DISPROVED__ or __UNKNOWN__ because of objective factors rather than incapacity. You need to output the conclusion." input: {Facts} +
"The Conclusion is: " + {Conclusion}
output: {nl_solution}
Configuration Parameters: Model: Qwen2.5-7B-Instruct or LLaMA-3.1-8B-Instruct Template: qwen or llama3 Method: LoRA fine-tuning Learning Rate: 5.0e-6 Epochs: 4.0 Batch Size: 1 (per device) Gradient Accumulation: 4 steps Max Length: 1024-2048 tokens Scheduler: Cosine with no warmup Precision: FP16
Example: Fact-based Reasoning Task Instruction: You are a helpful assistant. You are given a conclusion and a set of facts. You need to determine if the conclusion is __PROVED__, __DISPROVED__ or __UNKNOWN__ because of objective factors rather than incapacity. You need to output the conclusion. Facts: The standby is not a keen but it is affirmative. That this transalpine is not ascosporic is valid. That this transalpine is a keen holds. This transalpine is not a feud if that this stephanion is both not a keen and a chemical does not hold. This transalpine is lymphatic. That this transalpine is not a keen but it is a feud does not stand. If this stephanion is not lymphatic but a feud then that it is a chemical is valid. Something is apiarian if it is a chemical. Somebody is not non-lymphatic if it is a keen. This transalpine is not a kind of a chemical if this stephanion is a keen. That this stephanion is both not Mycenaean and a feud does not stand. This transalpine is not lymphatic if this stephanion is a keen. If this transalpine is a feud this stephanion is not a keen. If this transalpine is a feud then this stephanion is not apiarian. This transalpine is not a keen if that this stephanion is a kind of non-lymphatic thing that is a feud does not hold. That something is not non-Bahraini and it is a Picea does not stand if it does not naturalize. This stephanion is a feud if this stephanion is a keen and it is lymphatic. Conclusion: This stephanion is apiarian. LLM Reasoning Output: Step1: void -> assump1: Let’s assume that that this stephanion is both not lymphatic and a feud is false. Step2: fact15 & assump1 -> int1: This transalpine is not a keen. Step3: int1 & fact3 -> int2: This is contradiction. Step4: [assump1] & int2 -> int3: This stephanion is not lymphatic but a feud. Step5: int3 & fact7 -> int4: This stephanion is a chemical. Step6: int4 & fact8 -> hypothesis: Thus the hypothesis is __PROVED__. LLM Output: __PROVED__ True Answer: __PROVED__
Example: Story-based Reasoning Task Instruction: Here are some examples: Example 1: Question: Which fish tank has more thermal energy? Facts: Matter is made of tiny particles called atoms. Atoms are always moving. The energy of moving atoms is called thermal energy. The total amount of thermal energy in matter depends on three things: the type of matter, the amount of matter, and how fast the atoms are moving. Temperature measures how hot or cold matter is. If the atoms in matter slow down, the temperature goes down. The matter now has both less thermal energy and a lower temperature. What happens if the amount of matter changes? A 1-kilogram brick at 70°F has half as much thermal energy as a 2-kilogram brick at 70°F. The two bricks have the same temperature, but the smaller brick has half as many atoms. So, it has half as much thermal energy. Answer: the colder fish tank Analysis: The facts clearly explain that thermal energy depends on the amount of matter and temperature. The colder fish tank has less thermal energy because temperature is lower. This proves the answer is correct. Output: __PROVED__ Example 2: Question: What is the capital of France? Facts: Paris is a major city in Europe. France is a country in Europe. Answer: Paris Analysis: The facts mention Paris as a major city in Europe and France as a country in Europe, but they don’t explicitly state that Paris is the capital of France. The facts are insufficient to prove the answer. Output: __UNKNOWN__ Now evaluate this case: Facts: Matter is made of tiny particles called atoms. Atoms are always moving. The energy of moving atoms is called thermal energy. The total amount of thermal energy in matter depends on three things: the type of matter, the amount of matter, and how fast the atoms are moving. Temperature measures how hot or cold matter is. If the atoms in matter slow down, the temperature goes down. The matter now has both less thermal energy and a lower temperature. A 1-kilogram brick at 70°F has half as much thermal energy as a 2-kilogram brick at 70°F. The two bricks have the same temperature, but the smaller brick has half as many atoms. So, it has half as much thermal energy. Question: Which fish tank has more thermal energy? Choices: A. the warmer fish tank B. the colder fish tank Answer: the colder fish tank LLM Reasoning Output: The facts clearly explain that thermal energy depends on the amount of matter and temperature. The colder fish tank has less thermal energy because temperature is lower. This proves the answer is correct. LLM Output: __PROVED__ True Answer: __PROVED__
Example: Direct Evaluation Prompt for Trained Models Prompt Template: Here is the Hypothesis: {conclusion} Now These are the ’facts’: {facts} Please carefully evaluate the relationship between the facts and the hypothesis. Return UNKNOWN if the facts are insufficient to make a definitive conclusion. Return __PROVED__ only if the facts support the hypothesis. Return __DISPROVED__ only if the facts contradict the hypothesis. Output Conclusion: __PROVED__, __DISPROVED__, or __UNKNOWN__. Usage: This simplified prompt is used for evaluating fine-tuned models during inference, without examples or additional context.
LLM Judge Template: DPO sample selection for Atomic Reasoning DPO Instruction: You are an expert in reasoning. Given a logical reasoning problem, you need to select the response that demonstrates the most atomic, step-by-step reasoning process. Choose the response that breaks down the reasoning into the smallest possible logical steps, where each step establishes only one basic relationship. Problem: {problem_statement} Response A: {response_a} Response B: {response_b} Evaluation Criteria: Atomic Steps: Each reasoning step should establish only one basic logical relationship Granularity: Steps should be as fine-grained as possible Clarity: Each step should be clearly defined and understandable Completeness: All necessary steps should be included Correctness: Each step should be logically valid Output Format: Choose the response (A or B) that best demonstrates atomic reasoning with the most granular step-by-step breakdown.
DPO Training for Atomic Reasoning DPO Data Format: instruction: "You are an expert in reasoning. Given a logical reasoning problem, you need to select the response that demonstrates the most atomic, step-by-step reasoning process. Choose the response that breaks down the reasoning into the smallest possible logical steps, where each step establishes only one basic relationship." input: {problem_statement} output: {chosen_response} rejected: {rejected_response} Configuration Parameters: Model: Qwen2.5-7B-Instruct or LLaMA-3.1-8B-Instruct Template: qwen or llama3 Method: DPO Learning Rate: 5.0e-7 Epochs: 3 Batch Size: 1 (per device) Gradient Accumulation: 4 steps Cut_off Length: 2048 tokens
DPO Testing Case 1: (Fact-based Dataset Reasoning) Problem: Given the following facts, determine if the conclusion is __PROVED__, __DISPROVED__, or __UNKNOWN__. Facts: All birds have feathers. Penguins are birds. Feathers provide insulation. If something provides insulation, it helps maintain body temperature. Conclusion: Penguins can maintain their body temperature. Response A (Atomic Reasoning): Step 1: Penguins are birds (given fact) Step 2: All birds have feathers (given fact) Step 3: Therefore, penguins have feathers (from Step 1 and Step 2) Step 4: Feathers provide insulation (given fact) Step 5: Therefore, penguins have insulation (from Step 3 and Step 4) Step 6: If something provides insulation, it helps maintain body temperature (given fact) Step 7: Therefore, penguins can maintain their body temperature (from Step 5 and Step 6) Response B (Non-Atomic Reasoning): Step 1: Penguins are birds and all birds have feathers, so penguins have feathers and feathers provide insulation, which means penguins have insulation and can maintain body temperature. DPO Choice: Response A (Atomic Reasoning) - Each step establishes only one basic relationship.
DPO Testing Case 2: (Fact-based Dataset Reasoning) Problem: Given the following facts, determine if the conclusion is __PROVED__, __DISPROVED__, or __UNKNOWN__. Facts: If it rains, the ground gets wet. If the ground gets wet, plants grow better. It is raining today. Plants need water to grow. Conclusion: Plants will grow better today. Response A (Atomic Reasoning): Step 1: It is raining today (given fact) Step 2: If it rains, the ground gets wet (given fact) Step 3: Therefore, the ground gets wet today (from Step 1 and Step 2) Step 4: If the ground gets wet, plants grow better (given fact) Step 5: Therefore, plants will grow better today (from Step 3 and Step 4) Response B (Non-Atomic Reasoning): Step 1: It’s raining today, so the ground gets wet, and since wet ground makes plants grow better, plants will grow better today. DPO Choice: Response A (Atomic Reasoning) - Each step establishes only one basic relationship.
DPO Testing Case 3: (Story-based Dataset Reasoning) Problem: Which object has more potential energy? Facts: Potential energy depends on height and mass. Object A has mass 2kg and height 5m. Object B has mass 1kg and height 10m. The formula is PE = mgh. Choices: A. Object A B. Object B Response A (Atomic Reasoning): Step 1: Object A has mass 2kg and height 5m (given fact) Step 2: Object B has mass 1kg and height 10m (given fact) Step 3: The formula is PE = mgh (given fact) Step 4: For Object A: PE = 2 × g × 5 = 10g Step 5: For Object B: PE = 1 × g × 10 = 10g Step 6: Both objects have the same potential energy Step 7: Therefore, neither has more potential energy Response B (Non-Atomic Reasoning): Step 1: Object A has 2kg at 5m and Object B has 1kg at 10m, so using PE = mgh, A has 10g and B has 10g, so they’re equal. DPO Choice: Response A (Atomic Reasoning) - Each step establishes only one basic relationship.
DPO Testing Case 4: (Story-based Dataset Reasoning) Problem: What happens when you mix an acid with a base? Facts: Acids have pH less than 7. Bases have pH greater than 7. When acid and base mix, they neutralize. Neutralization produces salt and water. The resulting solution has pH close to 7. Choices: A. The solution becomes more acidic B. The solution becomes neutral C. The solution becomes more basic Response A (Atomic Reasoning): Step 1: Acids have pH less than 7 (given fact) Step 2: Bases have pH greater than 7 (given fact) Step 3: When acid and base mix, they neutralize (given fact) Step 4: Neutralization produces salt and water (given fact) Step 5: The resulting solution has pH close to 7 (given fact) Step 6: pH close to 7 means the solution is neutral Step 7: Therefore, the solution becomes neutral Response B (Non-Atomic Reasoning): Step 1: Acids and bases neutralize when mixed, producing salt and water with pH close to 7, so the solution becomes neutral. DPO Choice: Response A (Atomic Reasoning) - Each step establishes only one basic relationship.